OpenStudy (anonymous):

X/2 -10=-10 solving two step equation show work

OpenStudy (annon):

x/2=-10+10 x=0/2 x=0

OpenStudy (anonymous):

x=o because 0/2=0 and that minus -10 =-10

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

welcome